What are the steps to the problem ∛(a^2 b^2 )∙√(a^2 b^3 )? I already have the solution: ab^2)∙6√(a^4∙b). Thank you!
Answer by lwsshak3(11628)

What are the steps to the problem ∛(a^2 b^2 )∙√(a^2 b^3 )? I already have the solution: ab^2)∙6√(a^4∙b).
**
key notation:
^3√ cube root symbol
√ square root symbol
^6√ 6th root symbol
..
^3√(a^2*b^2)*√(a^2*b^3)
change to common 6th root (1/3=2/6, 1/2=3/6)
^6√(a^4*b^4)*^6√(a^6*b^9)
^6√(a^10*b^13)
(ab^2)^6√(a^4b)